When Sheila asked Margie to deliever camp supplies across the lake, they agreed to meet on the water. Shelia's boat can travel at 10km/h, while Margie's boat can do 12km/h. If the lake is 6.6 km wide and Sheila and Margie leave at the same time, how much time will it be before they meet?
Answer by funmath(2873)

Distance = rate*time or d=rt
They both travel the same amount of time:t (in hours)
Sheila's boat travels: 10t
Margie's boat travels: 12t
Altogether they travel: 6.6
So the problem we have to solve is:
10t+12t=6.6
(10+12)t=6.6
22t=6.6
22t/22=6.6/22
t=.3 hr
Most of us don't know off of the top of our head what .3 hrs is, so we might want to convert that to minutes.
t=.3(60 min)
t=18 min